# Comparing Hyper-optimized Machine Learning Models for Predicting Efficiency Degradation in Organic Solar Cells

### v2

@@ 2.3 Regression Problem
After observing in Section 2.2 the behavior of PCE over time for the various OSCs contained in our database, it was validated that the LMS regression fitting models yielding the best results were those with non-linear characterization. Table 2 displays the selected models along with their expressions as a function of time, dependent on the adjustment coefficients. It should be noted that the capability of these classical models lies solely in modeling the univariate behavior of the time effect on the PCE values. While it is possible to explore other LMS regression fitting in the multivariate domain, they only allow for establishing linear relationships, which do not adequately capture the behavior of our devices.
Table 2: LMS Bayesian regression fitting models to estimate PCE, denoted as $f(x)$, where $x$ represents time by means of parametric models.
@@
2.4 Machine Learning framework
In contrast to the previous classical approaches, ML moves forward to produce non-parametric regression models that adjust more complex behaviors without specifying the relationship between the output and input data by certain predetermined regression function. They may simply predict responses with new test data using a trained model. In this work, we exploit the advantages of a software framework, ROBERT ROBERT , developed under Python that facilitates hyper-optimization and benchmarking over well-recognized ML regression models by single command line instruction. This framework consists of the following modules:
  • Data curation: It processes the target dataset in order to filter correlated descriptors, noise, duplicates, as well as to identify and to convert categorical variables into one-hot descriptors.
  • Model selection: It iterates through multiple hyper-optimized models from scikit-learn pedregosa11a . It includes training size analysis by varying data partitions percentages (training-validation) and it also examines different data splitting options like k-neighbors clustering-based (KN) and random (RND). Permutation feature importance (PFI) detection is also performed. In consequence, models are computed with PFI-filtered descriptors and without them. As a default setting, this threshold eliminates features whose contribution to the model’s R2 is less than 4%. Moreover, it provides with insights on the dependencies of the input variables, thanks to SHapley Additive exPlanations (SHAP) analysis. Besides this, heatmap graphics are presented in order to compare models in terms of RMSE error. Amongst the generated models, it assesses the performance of Random Forests (RF), Multivariate Linear Models (MVL), Gradient Boosting (GB), Gaussian Process (GP), AdaBoost Regressor (AdaB), MLP Regressor Neural Network (NN), and Voting Regressor (VR) pedregosa11a .
  • Prediction: Once the models are trained and validated, the framework permits external test sets to predict target values with the best models.
  • Verify module: It assesses the predictive ability of the models, considering tests such as y-shuffle, y-mean, k-fold cross-validation, and one-hot features prediction.
  • Report module: With the aim of providing reproducibility and transparency, this module offers a detailed report containing comprehensive information about the ML models utilized and replication instructions through command line executions.
As an example of performance, Figure 4 illustrates the result obtained by ROBERT with our dataset as input data. Specifically, Figure 4 and Figure 4 compare the evolution exhibited by the ML models over time. The first figure depicts the result of a single model for training data up to 30 days, while the second figure does the same for training data up to 180 days. Their validity can be checked by inspecting the predicted outcome against the target values used as training (blue) and validation (yellow) data.
Figure 4: Results of the best ML models, tested with training (blue) and validation (yellow) target data up to: (a) 30 days and (b) 180 days.
@@ 3 Results
This section introduces the results obtained using both classical LMS regression fitting models and the optimal ML models as generated by the framework used in this work, to estimate the temporal behavior of our OSCs. The selected error metrics for analysis are briefly presented below:
  • Coefficient of determination R2: It quantifies how well the independent variables explain the variability of the dependent variable. Higher values indicate that the model fits the data well and captures larger proportion of the variability in the dependent variable.
@@
  • MAE: Another common objective error that measures the average magnitude of the errors between the predicted values and the actual values of the target variable. Unlike RMSE, which penalizes large errors more heavily, MAE treats all errors equally by taking the average of their absolute values. This makes MAE more robust to outliers in the data compared to RMSE.
@@
Meanwhile, SSE was defined in (4) and SStot is the total sum of squares, whereas $n$ is the number of samples, $y_{i}$ is the real observed value and $\hat{y}_{i}$ is the predicted value.
3.1 Classical LMS regression fitting results
@@
Hence, Figure 5 presents the accuracy metrics of the five Bayesian regression models introduced in Table 2 to estimate the performance of PCE over 180 days, for each of the OSCs available in the database (Table 1). It is worth noting that these results correspond to the mean values, along with the standard deviation for each OSC. Specifically, the values of R2 are shown in Figure 5, RMSE in Figure 5, SSE in Figure 5, and MAE in Figure 5. It can be observed that these models perform effectively when the temporal range of the fitting data is limited: as long as the time goes by, all models experience an increase in error, while R2 decreases. According to the results of Figure 5, it is observed that on average, the $gauss2$ model, consisting of two Gaussian terms, is the one that best estimates the performance of the OSCs in terms of PCE over time. However, it should be noted that for a 30 day fit, this model requires more points to achieve a valid R2. On the other hand, it should be pointed out that the maximum errors for all temporal fittings are bounded to values $\sim$0.06, fact that represents an error $>$1% of the the target value, the PCE. Notwithstanding that, it must be reiterated the limitation of these methods to characterize the behavior of all devices contained in our database in a global manner. These results reflect solely the mean performance of fitting for single OSCs, in turn.
@@
Figure 5: Accuracy metrics of the LMS-supported Bayesian regression fitting over time. Five different models are evaluated: $exp1$blacksquare, $exp2$blacksquare, $gauss1$blacksquare, $gauss2$blacksquare, $poly3$blacksquare. (a) Coefficient of determination, R2. (b) Root mean squared error, RMSE. (c) Sum of squared errors, SSE. (d) Mean absolute error, MAE.
@@
Figure 6: PCE estimation over time with the best LMS-supported Bayesian regression fitting model ($gauss2$). Four temporal datasets are used to compute the fittings: 30 days $\circ$-- ∘ -; 60 days $\circ$-- ∘ -; 90 days $\circ$-- ∘ - and 120 days $\circ$-- ∘ -.
3.1.1 PCE fitting estimation
Once preliminary fitting have been made with traditional LMS regression methods, it has been found that the best fitting corresponds to a double-term Gaussian function. Now we analyze the robustness of this method to estimate the behavior of PCE against the temporal variable. For this purpose, different estimations have been obtained, as shown in Figure 6. This figure presents estimated PCE data from a specific OSC device, namely $Cell4$ in our databased, which has been used so far in the all the previous examples throughout this work. Here, the parametric model is obtained with fitting data at: 30, 60, 90, and 120 days, respectively. Then the behavior of PCE at 180 days has been estimated, in order to be tested against the real data of the database. It can be observed that the estimated trend for fitting models with 30, 60, and 90 days deviate considerably from the real behavior (represented by the black line). It can be concluded that fitting data up to at least 120 days are required to make a robust fitting that achieves sufficient confidence in estimating future data at 180 days. This fact demonstrates that even the best parametric LMS fitting regression model requires more than half of the temporal data to reliably model the PCE evolution at future time values. In order to get further insights on the estimations plotted in Figure 6 more easily, Table 3 presents the specific details of these estimations according to the error metrics RMSE, SSE, and MAE. Upon inspection, the perception extracted from the previous figure is validated: the errors are unacceptable for any metric unless the fit is made with more than half of the data (90 days), and only if the subsequent PCE estimation is tested over time values immediately posterior to the temporal limit of the fitting.
@@ 3.1.1 PCE fitting estimation
Table 3: Accuracy metrics of the best LMS-supported Bayesian regression fitting model ($gauss2$) to estimate PCE over time, presented in Fig 6.
Fitting data Estimation RMSE SSE MAE
@@
3.2 Machine Learning regression results
In this section, we conduct the generation of ML models to learn from the temporal behavior of PCE versus time. It is noteworthy that, unlike the previous LMS-based regression fittings presented in Section 3.1, these ML models allow us to characterize the performance of our devices in a multivariate manner, thereby encoding the variability presented in their manufacturing, as well as in the particular environmental conditions at each temporal measurement, as indicated in Table 2.
Thanks to the use of the tool ROBERT ROBERT , we achieve an exploration of various combinations of ML algorithms and partition sizes. In order to maintain the same accuracy estimation policy used in Section 3.1, Figure 7 presents the same metrics as Figure 5. In this case, for each temporal training dataset (X-axis), the coefficient R2 is presented in Figure 7, RMSE in Figure 7, SSE in Figure 7, and MAE in Figure 7, for the best ML model in each case. It can be observed that, in general, all errors are bounded approximately within ranges $\sim$[0.02-0.03], for training data of 120 days onwards. This implies an error percentage lower than 1%, with respect to the mean value of PCE for these OSCs, which is highly satisfactory for modeling. It is also worth noting the good performance of the RF model, with training-validation partitions of 90-10%.
So as to comparatively assess the remaining ML models, Tables 45678, and 9, respectively present detailed accuracy results for the other models (GB, NN, MVL, and RF). Results with different training-validation ratios are presented: from 90-10% to 60-40%; while training data range from 30 to 180 days. For each table, the robustest model is highlighted in bold, considering the overall performance across all presented metrics, with priority given to the R2 value. It is worth mentioning that, even though they are not directly comparable models, it is confirmed that the errors of the ML methods are clearly bounded, while the LMS-based fitting regression errors grow with time. Furthermore, ML models have the ability to globally model a multivariate database, in contrast to univariate Bayesian fitting, which is independent for each device.
Figure 7: Accuracy of the best Machine Learning regression models over time: 30 days RF-60-10 $\blacksquare$; 60 days RF-90-10 $\blacksquare$; 90 days RF-90-10 $\blacksquare$; 120 days RF-90-10 $\blacksquare$; 150 days RF-90-10 $\blacksquare$ and 180 days RB-90-10 $\blacksquare$. (a) Coefficient of determination, R2. (b) Root mean squared error, RMSE. (c) Sum of squared errors, SSE. (d) Mean absolute error, MAE.
@@ 3.2 Machine Learning regression results
Table 4: Accuracy metrics of the Machine Learning regression models computed with training data up to 30 days.

### v3

@@ 2.3 Regression Problem with Classical Approaches
After observing in Section 2.2 the behavior of PCE over time for the various OSCs contained in our database, it suggests that the LS regression fitting models yielding the best results are those with non-linear characterization. Table 2 displays the selected models along with their expressions as a function of time, dependent on the adjustment coefficients. It should be noted that the capability of these classical models lies solely in modeling the univariate behavior of the time effect on the PCE values. While it is possible to explore other LS regression fitting in the multivariate domain, they only allow for establishing linear relationships, which do not adequately capture the behavior of our devices.
Table 2: LS Bayesian regression fitting models to estimate PCE, denoted as $f(x)$, where $x$ represents time by means of parametric models.
@@
2.4 ML framework
In contrast to the previous classical approaches, ML moves forward to produce non-parametric regression models that adjust more complex behaviors. In this work, we exploit the advantages of a software framework developed under Python, ROBERT ROBERT (see C for reproducibility details), that facilitates hyper-optimization and benchmarking over well-recognized ML regression models by single command line instruction. This automated framework consists of the following modules (for further details please check the online documentation Readthedocs ):
  • CURATE (data curation): It processes the input dataset in order to filter correlated descriptors, noise, duplicates, as well as to identify and to convert categorical variables into one-hot descriptors.
  • GENERATE (model selection): It iterates through multiple hyper-optimized models from scikit-learn pedregosa11a , including Random Forest (RF), Multivariate Linear Model (MVL), Gradient Boosting (GB), Gaussian Process (GP), AdaBoost Regressor (AdaB), MLP Regressor Neural Network (NN), and Voting Regressor (VR). The algorithms are combined with different training-validation split sizes, from 60-40% to 90-10% using random data splitting. For each combination of algorithm and training size, two models are generated: one with all the descriptors and another with only the most important variables detected by permutation feature importance (PFI) analysis. Among all the possibilities, the program selects two optimal models based on RMSE error.
  • PREDICT (external predictions selection): the framework is able to predict new target values. Moreover, it provides feature importances using SHapley Additive exPlanations (SHAP) and Permutance Feature Importance (PFI) analysis, and outlier detection.
  • VERIFY (assesing predictive ability): It assesses the predictive ability of the models, considering tests such as y-shuffle, y-mean, k-fold cross-validation, and prediction with one-hot encoding.
  • REPORT (generation of PDF reports): With the aim to enhance reproducibility and transparency, this module offers a detailed report containing comprehensive information about the ML models utilized and replication instructions through command line executions.
@@ 3 Results
This section introduces the results obtained using both classical LS regression fitting and ML models to estimate the temporal behavior of our OSCs. The selected error metrics for analysis are briefly presented below:
  • Coefficient of determination R2: It quantifies how well the independent variables explain the variability of the dependent variable. Higher values indicate that the model fits the data well and captures a larger proportion of the variability in the dependent variable.
@@
    where $n$ is the number of samples, $y_{i}$ is the real observed value and $\hat{y}_{i}$ is the predicted value.
  • MAE: Another common objective error that measures the average magnitude of the errors between the predicted values and the actual values of the target variable. Unlike RMSE, which penalizes large errors more heavily, MAE treats all errors equally by taking the average of their absolute values.
@@
3.1 Classical LS regression fitting results
Figure 4 presents the accuracy metrics of the five Bayesian regression models introduced in Table 2 to estimate the performance of PCE over 180 days, for each of the OSCs available in the database (Table 1). It is worth noting that these results correspond to the mean values, along with the standard deviation for each OSC. In general terms, it can be observed that these models tend to perform slightly better over short time periods compared to the longer experiments.
@@ 3.1 Classical LS regression fitting results
The maximum errors for all temporal fittings are bounded to values $\sim$0.06. According to the results of Figure 4, it is observed that on average, the $gauss2$ model, consisting of two Gaussian terms, is the one that best estimates the performance of the OSCs in terms of PCE over time. However, it should be noted that for a 30-day fit, this model requires more points to achieve a valid R2. It is worth mentioning that these methods cannot characterize the behavior of all devices contained in our database in a global manner, and the results obtained reflect solely the mean performance of fitting for individual OSCs.
@@
Figure 4: Accuracy metrics of the LS-supported Bayesian regression fitting over time. Five different models are evaluated: $exp1$blacksquare, $exp2$blacksquare, $gauss1$blacksquare, $gauss2$blacksquare, $poly3$blacksquare. (a) Coefficient of determination, R2. (b) Root mean squared error, RMSE. (c) Sum of squared errors, SSE. (d) Mean absolute error, MAE.
3.1.1 PCE prediction with LS
Once preliminary fittings have been made with traditional LS regression methods, it has been found that the best fitting corresponds to a double-term Gaussian function. Now we analyze the robustness of this method to predict the behavior of PCE against the temporal variable. For this purpose, four regression fittings have been obtained with PCE data acquired up to: 30, 60, 90, and 120 days, respectively. Then the behavior of PCE at 180 days has been predicted for each one (Figure 5 and Table 3). It can be confirmed that only the fitting with data up to 120 days is able to predict reliably future PCE values. This result demonstrates that even the best parametric LS fitting regression model requires more than half of the temporal data to reliably model the PCE evolution at future time values.
@@ 3.1.1 PCE prediction with LS
Figure 5: PCE predition over time with the best LS-supported Bayesian regression fitting model ($gauss2$). Four temporal datasets are used to compute the fittings: 30 days $\circ$-- ∘ -; 60 days $\circ$-- ∘ -; 90 days $\circ$-- ∘ - and 120 days $\circ$-- ∘ -.
@@
Table 3: Accuracy metrics of the best LS-supported Bayesian regression fitting model ($gauss2$) to predict PCE over time, presented in Figure 5.
Fitting data Prediction RMSE SSE MAE
@@
3.2 ML regression results
@@
Table 4: Accuracy metrics of the ML regression models computed with training data up to 180 days.
Training data ML model R2 RMSE SSE MAE
180 days GB-90-10 0.96 0.0640 0.0595 0.0520
GB-80-20 0.96 0.0720 0.0600 0.0560
GB-70-30 0.94 0.0840 0.0672 0.0650
GB-60-40 0.95 0.0700 0.0431 0.0500
NN-90-10 0.66 0.1400 0.3732 0.1100
NN-80-20 0.76 0.1100 0.1138 0.0670
NN-70-30 0.60 0.1600 0.3555 0.1200
NN-60-40 0.43 0.2500 0.7265 0.1800
MVL-90-10 0.81 0.1000 0.2929 0.0810
MVL-80-20 0.81 0.1000 0.2783 0.0800
MVL-70-30 0.74 0.1400 0.2696 0.1000
MVL-60-40 0.74 0.1300 0.2802 0.0920
RF-90-10 0.96 0.0480 0.0246 0.0390
RF-80-20 0.97 0.0520 0.0309 0.0370
RF-70-30 0.94 0.0670 0.0398 0.0460
RF-60-40 0.93 0.0850 0.0644 0.0580
In this section, we generate ML models to study the temporal behavior of PCE. Notice that, unlike the previous LS-based regression fittings presented in Section 3.1, these ML models allow us to characterize the performance of all the OSC devices under a single model, while also using the multiple variables in the dataset. These include manufacturing parameters and environmental conditions at each measurement. To comparatively assess the ML models, Table 4 presents detailed accuracy results for the algorithms GB, NN, MVL, and RF, with training data up to 180 days. Please note that all these models have PFI enabled, so that the most important features are considered. For further details about these features see Section 3.2.2 and C.
The nomenclature for the models is: ALGORITHM-TRAINING-VALID (e.g. RF-90-10). Results with different training-validation ratios are presented: from 90-10% to 60-40%. In this study, the optimal ML models used RF algorithms with a 90-10 training-validation partitioning. It can be observed that, in general, all MAE and RMSE values are bounded within ranges $\sim$[0.02-0.03], for training data of 120 days onwards (Figure 6).
The most robust model is highlighted in bold, considering the overall performance across all presented metrics, with priority given to the R2 value. Please note that the supplementary material (A) contains these same metrics for further inspection of models obtained with other temporal ranges. It is worth noting that errors of the ML methods are clearly bounded, regardless they operate on a multivariate database, in contrast to univariate Bayesian fitting presented in Section 3.1.
Figure 6: Accuracy of the best ML regression models over time: 30 days RF-60-40 $\blacksquare$; 60 days RF-90-10 $\blacksquare$; 90 days RF-90-10 $\blacksquare$; 120 days RF-90-10 $\blacksquare$; 150 days RF-90-10 $\blacksquare$ and 180 days RB-90-10 $\blacksquare$. (a) Coefficient of determination, R2. (b) Root mean squared error, RMSE. (c) Sum of squared errors, SSE. (d) Mean absolute error, MAE.
3.2.1 PCE prediction with ML
After comparatively evaluating the performance of ML models, this section validates their ability to predict the temporal behavior of the PCE for an OSC device never seen by the models. Firstly, the models are trained and validated without data of $Cell4$. Standing out above the others are the GB-90-10 and RF-90-10 models. In line with previous results (Table 4), and thus considering its robustness to model the entire dataset, RF-90-10 has been selected to get predictive inferences up to 180 days for the unseen OSC $Cell4$. Figure 7 provides such results. In particular, Figure 7 presents the results of the model when the dataset is splitted into traning and validation, at 90%-10% (blue and yellow points), whereas Figure 7 shows the prediction for $Cell4$ when its data are used as external test.
Next, Figure 8 presents validation tests. Figure 8 shows the RMSE error of the model compared to those of the tests: y-mean, y-shuffle and onehot. Figure 8 produces a 5-fold cross-validation test CV5F , with satisfactory results. In consequence, neither data leakage nor overfitting is evidenced.
Figure 7: Results of the best ML model, RF-90-10: (a) training-validation without data of $Cell4$ (R2=0.96, MAE=0.05, RMSE=0.071); (b) prediction of $Cell4$ as external test (R2=0.88, MAE=0.015, RMSE=0.021).
Figure 8: Validation of the best ML model, RF-90-10: (a) RMSE values for validation tests y-mean, y-shuffle and onehot; (b) 5-fold cross-validation test (R2=0.89, MAE=0.06, RMSE=0.09).
Finally, Figure 9 compares the temporal evolution of the PCE for $Cell4$ with the predicted data obtained by the model.
Figure 9: PCE prediction over time with the best ML regression model (RF-90-10, blue) for an OSC not seen during the training-validation process ($Cell4$, black).
3.2.2 Feature analysis
Considering the results obtained in the previous section, it is worthwhile to study the top-performing ML model that characterizes the behavior of OSCs (i.e. RF-90-10). ML models trained with a limited number of data points are unable to characterize and predict with sufficient accuracy. For this reason, we focused on analyzing the complete dataset, with measures up to 180 days. In this regard, PFI and SHAP analyses are presented below, by means of Figure 10 and Figure 11, respectively. PFI evaluates the significance of individual features in a ML model by measuring the increase in the model’s prediction error after permuting the values of a specific feature. The resulting increase in error indicates the dependency on that feature PFI1 ; PFI2 . Besides, SHAP permits interpreting individual predictions of ML models by computing the contribution of each feature to the model’s as an additive feature attribution method SHAP2017 .
Then Figure 10 presents the influence of the most relevant feature for the model, after applying PFI filtering, that is, removing variables with low effect on the R2. It is confirmed that the amount of solvent in the HTL layer, i.e. the amount of PEDOT:PSS, has the most significant influence. Likewise, the P3HT:PCBM ratio also demonstrates certain relevance. Additionally, dependencies with the value of PCBM are also observed, since it constraints the P3HT:PCBM ratio, given its non-linear influence in the denominator.
As for the environmental conditions, it is known that low humidity proves to be beneficial for these OSCs, which in our geographical location is normally correlated with high temperatures and atmospheric pressures. In a similar manner it acts the dew point, which directly correlates with humidity. Nonetheless, the device encapsulation demonstrates that the effect of these variables is minimized, as PFI filtering proves that they are not relevant enough to influence substantially the model.
Figure 10: PFI results for the best ML model generated with data up to 180 days.
Regarding the SHAP analysis, Figure 11 validates the previous insights: high values of the amount of solvent PEDOT:PSS in the HTL layer, high ratios of P3HT:PCBM are relevant to get stable PCE values. Besides this, the effect of the PCBM value on the P3HT:PCBM ratio is again confirmed.
Overall, the importance of PEDOT:PSS has been demonstrated as it plays an essential role in the multilayer structure of the OSCs. It confirms its relevance as the second layer to cover sufficiently the substrate. Moreover, the ratio P3HT:PCBM also demonstrates its role in dealing as charge carrier in the HTL layer.
Figure 11: SHAP results for the best ML model generated with data up to 180 days.
4 Discussion and conclusions
This paper has presented the application of optimal ML frameworks to characterize the degradation behavior, in terms of PCE, of OSCs with multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/Al. A dataset with 166 entries entries was created, containing PCE values of various OSC devices measured over 180 days. This dataset was supplemented with seven variables describing environmental conditions of the experiments and manufacturing parameters of the devices. Through hyper-optimization of a set of ML models we have presented an accuracy analysis of different methods, which were fed with OSC data periodized into sets from 30 to 180 days. The benchmarking has confirmed the validity of models like RF or GB to confer R2 values over 0.90, reaching in some extents R2$\sim$0.96-0.97 and error metrics (RMSE, SSE and MAE) significantly low when long term data is used for training. To reinforce the suitability of these ML models, classical LS regression fitting methods have been compared. These proved not to be suitable for a multivariable dataset like ours, especially when dealing with long term data of the OSCs. Consequently, their ability to predict PCE values is highly unreliable.
ML models proved to offer high feasibility to predict the behavior of unknown OSCs. Feature analysis suggests that the most influential variable is the solvent in the HTL layer, i.e., the amount of PEDOT:PSS. This is explained as the multilayer structure of the OSCs needs a minimum value of PEDOT:PSS to ensure a layer that completely covers the substrate. Moreover, the ratio P3HT:PCBM also exhibits significant importance, being higher values representative of greater impact on the model. Finally, it has been observed that variables such as temperature, humidity, dew point and pressure have lesser impact on the models, explained by the encapsulation made to the OSCs during their manufacturing.
CRediT authorship contribution statement
Conceptualization D.V. and J.V.A.-R.; Methodology D.V., J.V.A.-R., D.D. and F.R.-M.; Software D.V., J.V.A.-R., D.D. and M.F.; Data curation D.V. and M.F.; Formal analysis D.V., F.R.-M.; Funding acquisition D.V., J.C.F. and J.V.A.-R.; Investigation D.V., F.R.-M. and J.V.A.-R.; Resources F.R.-M. and J.C.F.; Supervision J.C.F. and J.V.A.-R.; Validation D.D. and M.F.; Visualization D.V. and M.F; Writing - Original draft D.V. and J.V.A.-R.; Writing - Review & Editing J.V.A.-R., D.D. and F.R.-M.
Acknowledgements
This work is part of the project CIAICO/2023/193 funded by Generalitat Valenciana. The authors acknowledge that support.
J.V.A.-R. and D.D. acknowledge Gobierno de Aragón Fondo Social Europeo (Research Group E07_23R), the State Research Agency of Spain (MCIN/ AEI/ 10.13039/501100011033/ FEDER, UE) for financial support (PID2022-140159NA-I00) and the European Union’s Recovery and Resilience Facility-Next Generation in the framework of the General Invitation of the Spanish Government’s public business entity Red.es to participate in talent attraction and retention programmes within Investment 4 of Component 19 of the Recovery, Transformation and Resilience Plan (MOMENTUM, MMT24-ISQCH-01).
Appendix A Extended accuracy metrics of ML models
This section comprises extended accuracy metrics of the ML models used in this work, when the temporal scope of the training-validation data is varied: 30, 60, 90, 120 and 150 days, respectively.
Table 5: Accuracy metrics of the ML regression models computed with training data up to 30 days.